# Emergence of two inertial sub-ranges in solar wind turbulence: dependence on heliospheric distance and solar activity

### v1

@@ 4.2 Observations during Solar Maximum
We now perform a similar spectral and intermittency analysis using the set of intervals recorded during the solar maximum (see Table 1). While the previous section was confined to only analyzing FSW, in this section we take into consideration the three main solar wind types, namely FSW, SSW and the ASSW. Previous studies on spectra and intermittency mostly focused on FSW and SSW (Bruno et al., 2003; Di Mare et al., 2019; Carbone et al., 2021; Sorriso-Valvo et al., 2021). More recently, the spectral properties of ASSW, which exclusively permeates the heliosphere during periods of high solar activity, were also examined (D’Amicis et al., 2021, 2022). However, such studies did not include intermittency. Moreover, a comparative analysis between FSW, SSW and ASSW at solar maxima has not yet been conducted. Thus, in this section, we examine the intermittency properties of ASSW (Marsch et al., 1981; D’Amicis et al., 2011; D’Amicis & Bruno, 2015) in comparison with the other two types of wind using Ulysses data, during the ascending phase of solar cycle 23 (year 2001), at $R\simeq 1.5$ au.
In Fig. 8, we show examples of the magnetic field power spectral density for three intervals at solar maximum, one for each type of solar wind. Unlike the previous case during solar minima, in this case, there is no clear emergence of two sub-ranges, and the inertial range is consistently showing the typical Kolmogorov scaling, close to -5/3. In the fast stream only, it is possible to see a shallower range at low frequency, compatible with the usual $f^{-1}$ range. This is also not visible in the two slow wind types. Similar spectra were observed for all the other intervals (not shown).
@@
The variation of $K$ (defined in Section 3) as a function of $\tau$ is shown in Fig. 9 for all the intervals of FSW, ASSW and SSW listed in Table 1. Similar to what has been observed during solar minima, $K$ is found to be scale dependent, decreasing with the time scale $\tau$ and approaching the Gaussian value $K\simeq 1.67$ at $\tau>\sim 10^{4}$. This is again a clear indication of the non-universal nature of the distribution function of the magnetic field increments. However, just like for the spectra, in this case no clear break within the inertial range is evident for any of the three types of solar wind.
@@
The consolidated plot shown in Fig. 9 allows to perform a comparative study of intermittency among those three types of solar wind. As evident from the plots, turbulence in ASSW is moderately intermittent, characterized by a value of $K$ which is intermediate between that of the SSW with the strongest intermittency and that of the FSW having the weakest intermittency. Our observations are in agreement with the fact that, in the outer heliosphere, the SSW is in a state of more developed turbulence than ASSW and FSW. This can also be inferred from the broad inertial range in the magnetic power spectra exhibited by SSW extending to much lower frequencies compared to FSW (see Fig. 8), for which we observe a $f^{-1}$ break (Matthaeus & Goldstein, 1986; Chandran, 2018). Several studies have observed similar spectral characteristics for these two types of wind (Bruno & Carbone, 2013; D’Amicis et al., 2018, 2022). The power spectrum of ASSW shows similar nature as SSW, with a broad inertial range. However, it is to be noted that a recent study conducted by D’Amicis et al. (2018) did observe a $f^{-1}$ break in the spectra of ASSW at 1 au, which hints how the turbulence develops in ASSW understood by the broadening of the inertial range as it evolves with $R$. While studies by D’Amicis & Bruno (2015); D’Amicis et al. (2018, 2021) explain the high Alfvénicity of ASSW as due to its generation from coronal hole boundaries based on its composition and micro-physics, our findings on ASSW with an intermediate state of turbulence hints that the low speed of ASSW may be due to the intermixing of FSW and SSW inside the heliosphere.
@@ 5 SUMMARY AND CONCLUSION
In this paper, we report the existence of two distinct sub-regimes for the inertial range in the magnetic power spectrum of solar wind turbulence within and beyond 1 au. Although a single inertial range spectral power law has been traditionally observed (Bruno & Carbone, 2013), a few studies have also identified variations in the spectral indices of the magnetic power spectrum (Wicks et al., 2011; Sioulas et al., 2023b) and in the scaling exponents of higher-order structure functions (Wu et al., 2022, 2023). Additionally, Sorriso-Valvo et al. (2023) observed a break ($\tau_{K}$) in the scaling of kurtosis ($K$) within FSW intervals, suggesting a possible connection between this break and the $f^{-1}$ break due to their similar behavior with $R$ as discussed by Davis et al. (2023). However, our findings show that $\tau_{K}$ in the kurtosis scaling closely coincides with the break ($f_{b}$) observed in magnetic spectra separating the two sub-regimes characterized by $f^{-3/2}$ and $f^{-5/3}$ spectral power laws in both the inner as well as the outer heliosphere (see Fig. 4). The appearance of a double power-law in both the magnetic power spectrum and kurtosis (or normalized fourth-order moments) supports the robustness of this break, indicating the existence of a previously unidentified characteristic scale within the inertial range. Whereas the most probable explanation for the $f^{-5/3}$ regime can be obtained by the isotropic Kolmogorov phenomenology or anisotropic mhd turbulence with a weak ${\bf v}$-${\bf b}$ alignment in non-Alfvénic solar wind, the $f^{-3/2}$ regime could be reasonably associated with the anisotropic spectra along the strong ${\bf v}$-${\bf b}$ alignment (Kolmogorov, 1941; Goldreich & Sridhar, 1995; Boldyrev, 2006). Note that, we consciously eliminate the possibility of a $-3/2$ spectra by Iroshnikov-Kraichnan phenomenology which is valid only for balanced MHD and cannot explain the emergence of $-3/2$ spectra when there is a strong ${\bf v}$-${\bf b}$ correlation.
A recent study by Zhao et al. (2024) provided evidence of a transition from a weak to a strong turbulence regime as one moves from larger to smaller scales. In our study, an inspection of the cross-helicity co-spectra revealed that the turbulence in FSW shifts from a highly imbalanced state ($|z^{+2}|\gg|z^{-2}|$, or vice-versa) at larger scales to a relatively balanced one ($|z^{+2}|\sim|z^{-2}|$) on moving towards the smaller scales (see Fig. 6). These observations may explain the broken power-law behavior of the spectrum and the kurtosis indicating a transition in the nature of turbulence as the cascade progresses towards the smaller scales.
@@
We have further investigated the dependence of the sub-inertial regime break ($\tau_{K}$) on the heliospheric distance ($R$) in comparison with the ion and correlation scales. Our findings indicate a power-law behavior for $l_{K}$ (Taylor transformed $\tau_{K}$, Taylor, 1938) with $R$, which upon normalization with the typical ion scales (e.g. the ion-inertial scale $d_{i}$ and the ion gyro-radius $\rho_{i}$) and the correlation scale ($L_{c}$) practically disappears in the inner heliosphere (see Fig. 7). Therefore, both the correlation scale and the characteristic ion scale appear to control the location of the break. Interestingly, though, $l_{K}$ appears to approach the correlation scale shifting towards larger scales as $R$ increases, resulting in the absence of the $f^{-3/2}$ regime at $3.16$ au. This observation could explain the transition of the inertial range magnetic spectral slope from $-3/2$ near the Sun to $-5/3$ farther away (Chen et al., 2020; Shi, C. et al., 2021; Sioulas et al., 2023a). Note that a residual power-law radial dependence of the break scale still persists in the outer heliosphere, possibly due to variations in the latitude at which the FSW streams were sampled. This residual behaviour of the normalized $l_{K}$ must be studied in depth in a future study as functions of the latitude and also the large-scale magnetic field angle, which determines the degree of anisotropy in the measured turbulence.
@@
Our analysis, extended to a period of year 2001 for different types of solar wind intervals, shows that the two inertial sub-ranges are no longer surviving during the period of high solar activity. Nevertheless, the study enables us to characterize the state of turbulence in the Alfvénic slow solar wind (Marsch et al., 1981; D’Amicis et al., 2011; D’Amicis & Bruno, 2015; D’Amicis et al., 2021, 2022), as compared to traditional fast and slow winds. We showed that during this period, when ASSW is found in abundance near the ecliptic plane, it is in an intermediate state of turbulence between those typical of fast and slow streams. This also gives us insights on the position of the break separating the integral ($f^{-1}$) and inertial ranges (Matthaeus & Goldstein, 1986; Chandran, 2018) in the case of ASSW. While D’Amicis et al. (2018) found the $f^{-1}$ break to be occurring at the same frequency for FSW as well as ASSW at 1 au, in our study it occurs at a much lower frequency for the ASSW compared to the FSW at distances greater than 1.5 au (see Fig. 8), thus suggesting a plausible explanation for the ‘slowness’ of the ASSW due to strong intermixing between the FSW and the SSW during the high solar activity.
@@ Appendix B Component-wise kurtosis of the magnetic field fluctuations in FSW intervals during solar minimum
In this appendix we show the kurtosis $K$ for each individual RTN magnetic field component, using four example intervals from both Helios and Ulysses database, at eighth different distances from the Sun. Different colors refer to the different intervals. Whenever present, a power law is shown as colored dashed line, and the corresponding scaling exponents are indicated in each panel. Two power laws can be identified in all of the Helios and most of the Ulysses intervals, with the exception of the radial component at 2.75 au and of all components at 3.16 au. The timescale $\tau_{K}$ of the break between the two power laws is indicated by a solid vertical grey line, while the dashed grey vertical lines indicate the location of the spectral break, $1/f_{b}$.

### v2

@@ 4.2 Observations during Solar Maximum
We now perform a similar spectral and intermittency analysis using the set of intervals recorded during the solar maximum (see Table 1). While the previous section was confined to only analyzing FSW, in this section, we take into consideration the three main solar wind types, namely FSW, SSW and ASSW. Previous studies on spectra and intermittency mostly focused on FSW and SSW (Bruno et al., 2003; Di Mare et al., 2019; Carbone et al., 2021; Sorriso-Valvo et al., 2021). More recently, the spectral properties of ASSW, which permeates the heliosphere during periods of high solar activity, were also examined (D’Amicis et al., 2021, 2022). However, such studies did not include intermittency. Moreover, a comparative analysis between FSW, SSW and ASSW at solar maxima has not yet been conducted. Thus, in this section, we examine the intermittency properties of ASSW (Marsch et al., 1981; D’Amicis et al., 2011; D’Amicis & Bruno, 2015) in comparison with the other two types of wind using Ulysses data, during the ascending phase of solar cycle 23 (year 2001), at $R\simeq 1.5$ au.
In Fig. 8, the smoothed magnetic power spectra are shown for all intervals listed in Table 1. Compensated spectra in the relevant regions are shown in the bottom panels, where the values of the normalized cross-helicity are also indicated for each stream. In contrast to the solar minima, where two regimes with $f^{-3/2}$ and $f^{-5/3}$ were systematically found, during solar maximum we find FSW intervals both with and without the $f^{-3/2}$ regime. This could be due to the fact that the break between the $f^{-3/2}$ and $f^{-5/3}$ regime has evolved to larger scales beyond the correlation scale of turbulence, which was measured to be $\sim 1250$ s $(8\times 10^{-4}$ Hz)))) for these intervals. A similar observation was made in the previous section for the interval at $R=3.16$ au. Note also that a possible role of latitudinal variation (which spans from $-56^{\circ}$ to $75^{\circ}$ with respect to the ecliptic, see Table 1), was not apparent but cannot be excluded. For the ASSW, the $f^{-3/2}$ regime is also evident in the spectra for most of the intervals, with the exception of one case with lower cross-helicity. However, the $f^{-1}$ regime is not apparent, as the correlation scale in this case is $\sim 5000$ s $(2\times 10^{-4}$ Hz)))). On the other hand, the spectra of the SSW intervals exhibit a broad $f^{-5/3}$ regime extending to much lower frequencies with the $f^{-1}$ and $f^{-3/2}$ regimes being absent.
@@
The variation of $K$ (defined in Section 3) as a function of $\tau$ is shown in Fig. 9 for all the aforementioned intervals. Similar to what has been observed during solar minima, $K$ is found to be scale dependent, decreasing with the time scale $\tau$ and approaching the Gaussian value $K\simeq 1.67$ at $\tau>\sim 10^{4}$ s (See Appendix A). This is again a clear indication of the non-universal nature of the distribution function of the magnetic field increments. From Fig. 9 (a), (b) & (c), it is evident that a steeper power law followed by a shallower one is commonly observed for FSW and ASSW, while for SSW the shape varies quite a lot and it is hard to determine distinct regimes. The two power law regimes existing for FSW and ASSW are fairly consistent with the existence of two distinct regimes in the spectra of these two types of wind (for more details see Section 4.1). Moreover, for many intervals, the break in the kurtosis roughly corresponds to that in the power spectra between the $f^{-5/3}$ and $f^{-3/2}$ regimes in both FSW and ASSW, implying a close relation between them. For the SSW, where distinct regimes in kurtosis are not easily observed, the spectra is found to exhibit a broad $f^{-5/3}$ scaling.
@@
The consolidated plot shown in Fig. 9 (d) allows to perform a comparative study of intermittency among those three types of solar wind. As evident from the plots, turbulence in ASSW is moderately intermittent, characterized by a value of $K$ which is intermediate between that of the SSW with the strongest intermittency and that of the FSW having the weakest intermittency. Our observations are in agreement with the fact that, in the outer heliosphere, the SSW is in a state of more developed turbulence. This can also be inferred from the broader inertial range in the magnetic power spectra exhibited by SSW, extending to lower frequencies compared to FSW and ASSW.
We note that a recent study conducted by D’Amicis et al. (2018) observes a $f^{-1}$ break in the spectra of ASSW at 1 au, clearly showing how the turbulence develops in ASSW by the broadening of the inertial range as it evolves with $R$. Studies by D’Amicis & Bruno (2015); D’Amicis et al. (2018, 2021) explain the high Alfvénicity of ASSW as due to its generation from coronal hole boundaries based on its composition and micro-physics.
@@ 5 SUMMARY AND CONCLUSION
In this paper, we report the existence of two distinct sub-regimes for the inertial range in the magnetic power spectrum of solar wind turbulence within and beyond 1 au. Although a single inertial range spectral power law has been traditionally observed (Bruno & Carbone, 2013), a few studies have also identified variations in the spectral indices of the magnetic power spectrum (Wicks et al., 2011; Sioulas et al., 2023b) and in the scaling exponents of higher-order structure functions (Wu et al., 2022, 2023; Sorriso-Valvo et al., 2023).
@@
In a series of fast, Alfvénic solar wind intervals measured during low solar activity, our findings show that a clear break in the kurtosis scaling closely coincides with the break observed in magnetic spectra separating the two sub-regimes characterized by $f^{-3/2}$ and $f^{-5/3}$ spectral power laws, both in the inner and outer heliosphere (see Fig. 4). The appearance of a double power-law in both the magnetic power spectrum and kurtosis supports the robustness of this break, indicating the existence of a previously unidentified characteristic scale within the inertial range. Whereas the most probable explanation for the $f^{-5/3}$ regime can be obtained by the isotropic Kolmogorov phenomenology or anisotropic MHD turbulence with a weak ${\bf v}$-${\bf b}$ alignment in a non-Alfvénic regime of solar wind fluctuations, the $f^{-3/2}$ regime could be reasonably associated with the anisotropic spectra along the strong ${\bf v}$-${\bf b}$ alignment (Kolmogorov, 1941; Goldreich & Sridhar, 1995; Boldyrev, 2006). Note that we consciously overlook the possibility of a $-3/2$ spectra by Iroshnikov-Kraichnan phenomenology (Kraichnan, 1965), which is only valid for balanced MHD and cannot explain the emergence of $-3/2$ spectra when there is a strong ${\bf v}$-${\bf b}$ correlation. A recent study by Zhao et al. (2024) provided evidence of a transition from a weak to a strong turbulence regime as one moves from larger to smaller scales. In our study, an inspection of the cross-helicity co-spectra revealed that the turbulence in FSW shifts from a highly imbalanced state ($|z^{+2}|\gg|z^{-2}|$, or vice-versa) at larger scales to a relatively balanced one ($|z^{+2}|\sim|z^{-2}|$) on moving towards the smaller scales (see Fig. 6). These observations may explain the broken power-law behaviour of the spectrum and the kurtosis indicating a transition in the nature of turbulence as the cascade progresses towards the smaller scales.
@@
We have further investigated the dependence of the sub-inertial regime break on the heliospheric distance, also in comparison with the ion and correlation scales. Our findings indicate a power-law behaviour for $l_{K}$ (Taylor transformed $\tau_{K}$, Taylor, 1938) with $R$, which upon normalization with the typical ion scales (e.g. the ion-inertial scale $d_{i}$ and the ion gyro-radius $\rho_{i}$) and the correlation scale ($L_{c}$) practically disappears in the inner heliosphere (see Fig. 7). Therefore, both the correlation scale and the characteristic ion scale could be controlling the location of the break. Interestingly, $l_{K}$ appears to approach the correlation scale shifting towards larger scales as $R$ increases, resulting in the absence of the $f^{-3/2}$ regime at $3.16$ au. This observation could explain the apparent transition of the inertial range magnetic spectral slope from $-3/2$ near the Sun to $-5/3$ farther away recently reported in the inner heliosphere (Chen et al., 2020; Shi, C. et al., 2021; Sioulas et al., 2023a). Note that a residual power-law radial dependence of the break scale still persists in the outer heliosphere, possibly due to variations in the latitude at which the FSW streams were sampled. This residual behaviour of the normalized $l_{K}$ must be studied in depth in a future study as functions of the latitude and also the large-scale magnetic field angle, which may play a role in the degree of anisotropy in the measured turbulence. Finally, we have performed a complementary preliminary analysis during high solar activity. Our results show similar but less robust features as described above for fast and slow solar wind, with single-scaling inertial ranges occasionally observed in FSW and ASSW streams. Due to the reduced sample used here, a possible role of latitudinal dependence cannot be ruled out.
@@
The variability of wind types at solar maxima enables us to additionally characterize the state of turbulence in the Alfvénic slow solar wind, as compared to traditional fast and slow winds. ASSW, which is found in abundance near the ecliptic plane, is in an intermediate state of turbulence between typical fast and slow streams, with double scaling in the spectra but less regular scaling laws of the kurtosis. It is interesting that we do not observe the $f^{-1}$ break in the ASSW spectra (Matthaeus & Goldstein, 1986; Chandran, 2018), suggesting a possibility that this is located at a frequency lower than those accessible with our samples. Interestingly, for the ecliptic solar wind at 1 au, the $f^{-1}$ break was found at the same frequency for both FSW and ASSW D’Amicis et al. (2018). In the current study, at distances greater than 1.5 au, similar $f^{-1}$ break in ASSW occurs at a much lower frequency compared to that for the FSW (see Fig. 8).
The observations described in this article clearly indicate the emergence of a characteristic scale within the inertial range of Alfvénic solar wind turbulence, likely separating two different dynamical regimes. The absence of double scaling in SSW at solar maximum seems to preliminarily suggest that Alfvénicity might play a crucial role in the occurrence of such break. However, a physical interpretation requires the disentanglement of several additional possible factors (e.g., geometry, expansion, latitude, turbulence amplitude), which is only possible through an extended statistical analysis, deferred to future works. These findings could be of broad relevance for heliospheric studies, possibly informing solar wind modeling (e.g., Cranmer, 2012; Usmanov et al., 2012; Zank et al., 2021; Chandran, 2021), affecting energetic particle transport (Pucci et al., 2016; Shukurov et al., 2017; Perri et al., 2021), and more generally constraining the energetics of the solar wind (Vasquez et al., 2007; Rivera et al., 2024).
@@ Appendix B Component-wise kurtosis of the magnetic field fluctuations in FSW intervals during solar minimum
In this appendix we show the kurtosis $K$ for each individual RTN magnetic field component, using four example intervals from both Helios and Ulysses database, at eight different distances from the Sun. Different colours refer to different intervals. Whenever present, a power law is shown as coloured dashed line, and the corresponding scaling exponents are indicated in each panel. Two power laws can be identified in all of the Helios and most of the Ulysses intervals, with the exception of the radial component at 2.75 au and of all components at 3.16 au. The timescale $\tau_{K}$ of the break between the two power laws is indicated by a solid vertical grey line, while the dashed grey vertical lines indicate the location of the spectral break, $1/f_{b}$.